Back to the NIKA2 main page

Daily Reports

Jan. 10th, Tuesday

15:15-21:00; Miguel, Angel: Technical time (T06-22) scheduled up to Wed. 04:00 (LT). Checking interference spikes in EMIR data after asking CETURSA IT team for switching off suspicious Wi-Fi systems of the lift booths close to the telescope. In principle, no interference spikes were detected after ad-hoc technical observations. Tau_225GHz is about 0.25 and stable. Switching to NIKA2 DAQ started at 16:50; bad KIDS: 61 (ar1), 9 (ar2), 71 (ar3). Noise traces (on-sky) of ar2, ar1 and ar3 are given below. A quick pointing session (11 sources at 70 < Az. < 381 deg and 27 < El. < 69 deg) for refining corrections was carried out from 20:20 to 21:50.

attachment:traces_ar1_onsky_20230110.png attachment:traces_ar2_onsky_20230110.png attachment:traces_ar3_onsky_20230110.png

21:00-03:00; Manuel, Angel: Opacity keeps stable (tau_225GHz < 0.25). Pointing session was finished. Resulting statistics: <pAz.> = -0.46 +/- 1.38 arcsec, <pEl.> = 6.25 +/- 0.64 arcsec. Continued with a calib_1scan on Uranus (20220110s245), a skydip (s246), and a calib_1scan on ALFORI (s249). Only the parameter P7 of the pointing model was corrected accordingly and checked before scan s257. Fairly stable after a brief opacity increase. Pointing and focus on Uranus. Beam map on Uranus (s258). Starting science observations: project 105-22; one rep. (35 min) on B213-C6; project 122-16 (GASTON), 2 reps. on B213E_N; project 192-16 (N2CLS) started: 1 rep. on COSMOS field.

Jan. 11th, Wednesday

03:00-09:00; Jean-François: clear sky, Tau_225 stable around 0.25, peak-to-peak fluctuations < 5Jy/b and rms=2-3Jy/b at 2mm all night. Project 192-16 on COSMOS for 3 hours (hit attenuation limit); 199-16 PSZ2G212 for 1.5 hours (i.e. 3rep) (hit attenuation limit); focus; 169-22 GJ526North 1 hour; focus; start 160-16 NGC4254 and pass over to Lara.

[AB: this morning Dave reported noticeable pressure oscillations (e.g. P101) apparently linked to the dilution circuit of the cryostat. T_MC has been below 175 mK until now. The problem has been tackled by Dave and Alessandro and was related to the heater in the instrument, after discarding that the pre-amplifiers act as 'heaters']

09:00-15:00; Lara: three runs done for NGC4254 until 10:00. Clear sky, tau almost stable around 0.25, peak to peak fluctuations < +/- 6 Jy/b, rms < 2 Jy/b. After this moved to MWC349 for pointing, focus, calib_1scan + another calib_1scan on NGC7027. Then I did a skydip. At 11 some checkings with NIKA2 and EMIR were done for possible wifi contaminations. 13:35: cryostat intervened by Dave John and Alessandro Monfardini. P5 stabilised and 2mm pre-amplifier restarted. DAQ restarted at 14:00: bad KIDs: 54 (ar1), 9 (ar2), 65 (ar3). At 14:15 observations start again with pointing and focus on 2251+158. Start 160-16 NGC7331 after pointing on 2234+282. First run with velocity = 60, then changed to 45 since the elevation was ~ 75. A large value of p2p and high elevation: change to 159-22.

15:00-21:00; Angel: Opacity about 0.25-0.30 at the beginning of the shift. Evidence of instabilities in pointing and focus scans. Only NGC1569 (159-22) and NGC7331 (160-16) after a while are suitable for observation at this time with current opacity levels. Unfortunately, pressure oscillations appeared again, but this time they are apparently related to PT #2. Dave, Martino, and Alessandro are working on this issue. Switching to EMIR backup programs at 17h (LT). NIKA2 window was closed and the cap was assured with Al tape by request. Continuing on backup project 125-22 (3 mm). NCS failure: VAC1 and VAC3 restarted; the first focus was canceled due to a problem with the AntennaMountDrive services. Resuming observations at 18:25 after the intervention of Joaquín and William. A total of 3 beta and 6 lambda maps (tile 3) on the G111 molecular cloud were obtained, and 3 beta maps (tile 4) were launched at the end of the shift.

SL: I think there's a small mistake about the oscillations hypothesis: PT #2 is connected to the 4K plate, if oscillations were due to it, there would be noise lines in the data spectra perfectly visible. PT #1 however is connected to the 30K stage, a touch on it could generate oscillations possibly not visible in data spectra.

The telescope resuming observations at sunset:

attachment:IMG_0518_mod_sm.JPG attachment:IMG_0536_mod_sm2.JPG

21:00-03:00; Damien: Opacity was quite unstable for the whole shift, but pwv was ok (2~6mm). 10 reps (12min each) on L1544 for project 104-22; 23h50 switched to project 120-22, 8 reps (14min each) on L1495A-NS with setup 2. Started last 2 reps on L1495A-NS then let Jean-François take over.

Jan. 12th, Thursday

03:00-09:00; Jean-François: EMIR obs for whole run because of NIKA2 cryo pb and unstable atmosphere: Observe project EMIR 162-22 (source NSA036285) - Start project EMIR 124-22

09:00-15:00; Lara: Good opacity (~0.23). Observing project 124-22 with EMIR because of some checkings on NIKA2 cryostat had to be done (scans #95 - #106). [AB: after increasing the speed of the turbo-molecular pump HP400 by Dave, pressures and temperatures stabilised close to the expectations]. 10:30 switch to NIKA2. DAQ restarted at 10:40 (LT); bad KIDS: 49 (ar1), 6 (ar2), 65 (ar3). Noise traces are given below. Pointing and focus on K3-50A. Pointing on MWC349 (FWHM ~ 17 arcsec) + calib_1scan on MWC349 and NGC7027 (p2p = 2 Jy/b and p2p = 4 Jy/b respectively, at 2mm). Skydip launched at 11:20. At 11:35 good opacity (0.2) and stable weather conditions, clear sky. Moving to 1418+546, do pointing and focus on 1633+382. 12:10 start with project 199-16 on source PSZ2-G086.9 (tau 1.7-1.8 and target elevation about 50).

attachment:traces_ar1_onsky_20230112.png attachment:traces_ar2_onsky_20230112.png attachment:traces_ar3_onsky_20230112.png

15:00-21:00; Stefano. I take over from Lara around 14 local time. Completed the last repetition on PSZ2-G086.9 (199-16), continued with pointing, focus, two reps on PSZ2-G046.1, then starts the difficult part of the day. It's hard to focus at El. 35-40 deg. Need a higher elevation. Finally! Then moved to 171-22 and performed 4 reps on PSZ2G085. Usual pointing and focus checks regularly performed. After this, calib_1scan on CRL618 (scan 20230112s202). Then switched to 122-16: 2 reps on B213-N. IMEGIN (160-16) follows: NGC337, 5 repetitions (added 2 reps to compensate for older reps taken under sub-optimal conditions), Source NGC337 is completed!

Life signs on the moon:

attachment:IMG_0661_mod_sm.JPG attachment:IMG_0605_mod_sm.JPG

21:00-03:00; Damien: Good opacity (~0.22). Completed 10 reps (9-10 min each) on NGC1569 (Project 159-22). 23h : pointing, focus, calib_1scan (scan 20230112s265), and beam map (scan 266) on Uranus. 23h30: switched to project 199-16, 5 reps (22 min each) on PSZ1G226; focus then let Jean-Francois take over.

Jan. 13th, Friday

03:00-09:00; Jean-François: Clear sky, A2 beam 18"x18" or better, opacity stable 0.20-0.25, smooth run. Project 192-16 : COSMOS (2 hours). Project 199-16 : 3 repeats on PSZ2G212 to finish it (1.5 hours). Project 160-16 on NGC4254 (1 hour) (still 1 hour to go). Project 169-22: 1 repeat on GJ526North (1hr) (15 min. overboard).

09:00-15:00; Lara: Clear sky, opacity 0.19, beam 17x18 (2mm). Project 199-16 on PSZ2G087 (started at 09:45, scan #86-#106). Stop at 11:40 for test heater cables of the NIKA2 cryostat (Dave and Santiago). DAQ restarted at 13:00 (LT); bad KIDS: 48 (ar1), 6 (ar2), 67 (ar3). Skydip until 13:17. Focus on 2251+158 and calib1scan on NGC7027 and CRL2688. 13:36 moved to project 160-16 (NGC7331).

15:00-21:00; Stefano: start with 171-22 and 7 repetitions on ACT0022. Follows 160-16 with 4 repetitions on NGC 7331 (90 minutes in total). Finally switched to 159-22 and started observing NGC1569.

It's Friday, gotta love the 30m:

attachment:IMG_0711_mod_sm.JPG attachment:IMG_0682_mod1_sm.JPG attachment:sunset_20230113_sm.JPG

21:00-03:00; Damien : 7 reps (10' each) on NGC1569 (159-22); 22h10 : pointing, focus, calib1scan and beammap on Uranus. Then switched to 122-16, 2 reps on B213E_N (~90 min total). 00h30: Started project 199-16, 3 reps on PSZ1G226 (22' each ) to finish it.

Jan. 14th, Saturday

03:00-09:00; Jean-François: Clear sky, RH=10%, A2 beam 18"x18" or better, opacity stable 0.20-0.15, smooth run. Project 192-16: COSMOS (2 hours). Project 199-16: 5 repeats on PSZ2G087 to finish it (2 hours). Project 169-22: 1 repeat on GJ526North (1hr).

09:00-15:00; Lara: Stable opacity 0.17-0.2 and atmosphere. Start with project 160-16 with NGC4254 (observing the last hour required on this source before it becomes too low in elevation). 10:20 start with project 199-16 (on PSZ2G081). 12:00 focus on 2251+158 and calib_1scan on NGC7027 and CRL2688. 12:40 moved to project 160-16 (NGC7331).

15:00-21:00; Stefano: I take over and - after NGC7331 (120-16) - I start with project 171-22 and perform 6 repetitions of ACT0022. Then 122-16 with 3 repetitions of B213-N. Two repetitions on NGC7331 (project 160-16) follow, for a total of 44 minutes... and the NGC7331 planned observations are completed. Finally, I switch to 159-22 and start observing NGC1569.

21:00-03:00; Damien: opacity unstable during the shift, start with 5 reps on NGC1569 (50 min total). Then switch to nikaw-22 for a calib1scan, skydip and a beammap on Uranus. 23h30 start project 122-16, 2 reps on B213E_N (90 min total). 1h30; switch to project 199-16, 1 rep on PLCKG227 (22 min). then did a focus and pointing before starting a rep on COSMOS (192-16).

Chasing the comet ZTF (C/2022 E3):

attachment:IMG_4329_mod_sm.JPG attachment:IMG_4613_16_mod3.JPG

Jan. 15th, Sunday

03:00-09:00; Jean-François: Observing conditions were excellent for the first 4 hours (A2 beam 18"x18" or better, opacity stable 0.15-0.2, and p2p better than +/- 3Jy/b) for project 192-16: COSMOS (3 rep. 3 hours) and project 199-16: PSZ2G081 (2 rep. 0.5hr). Then conditions were moderately fair (A2 beam 18"x18", opacity 0.20-0.4, and p2p between +/-5 and +/-10Jyb) for project 169-22: GJ526North (2 rep. 1.8hr).

09:00-15:00; Lara: Clear sky but some spikes in the opacity by taumeter (0.3-0.5). After monitoring the atmosphere using nikaw-22 account (p2p < 4 Jy/b and rms ~ 0.5 on 1502+106), continue on the project 169-22 (starting at 09:50; tau ~ 0.3). Ends at 11:00. Pointing focus and calib1scan on MWC349. Spikes spikes spikes in the taumeter &&%*#&$ and high clouds passing by. Might improve in the next 2 hours. Switching to temporary time filler. Peak to peak reached +/- 50 Jy on scans with constant elevation. EMIR backup projects have been checked but the targets are too close to the Sun position. Thus, continue with NIKA2, while monitoring the atmosphere instabilities. Pointing on 1641+399 (with nikaw-22). Peak to peak +/- 10-20 Jy on pointing. No projects accept these conditions. We stop and wait a while. Then we decide to switch to EMIR at the end of this shift.

[it is painful to see that E090 is observed while tau225~0.16 and p2p~9Jy/b @ 2mm for NIKA2 in scan 128.... <CK>]

15:00-21:00; Stefano. Lots of cirrus everywhere. I'm learning some 1-pixel mapping. Observing project 125-22. In fact, I start to enjoy it and every completed map I'm getting prouder of the result. The sky is getting more cloudy (15:30 UT), both above our heads and under our feet. The tau-meter tau stays low at 0.16. Grrrrr! @16:00 UT the taumeter has realised that we're immersed in the clouds and jumps to 0.3. LOL. At roughly 16:30 UT we stop the telescope, overwhelmed by the 20mm pwv.

It was a nice and sunny day at the 30m telescope...

Special time filler observations:


21:00-03:00; Damien: Switching to NIKA2 DAQ started at 22:00; bad KIDS: 42 (ar1), 6 (ar2), 66 (ar3). Stable opacity (0.17~0.2)> Pointing, focus, calib_1scan (scan #180), and beam map (scan #181) on Uranus. Then switch to project 122-16, 1 rep to finish B213E_N 1st script (44 min) and 1 rep on B213E_N 2nd script (44 min). Switch to project 199-16, 3 reps on PLCKG227 (66 min total). Finally, skydip then start a rep on COSMOS (192-16).

Jan. 16th, Monday

03:00-09:00; Jean-François: during the run, a technical problem occurred: lost resonances on A2; Angel had to be awakened; he had done a hard reset of the electronic crate of Array2 in the cabin according to a procedure devised by Dave; the crate restarted at 5:15 local; numbers of lost kids after reset: 53, 7, 70 on A1, A2, A3, respectively; in total one hour lost. Observing conditions were fairly good for the first 2 hours (A2 beam 18"x18" or better, opacity stable around 0.2, and p2p better than +/- 6Jy/b) when project 192-16 COSMOS was observed (2 repeats, 2 hours). Then conditions were moderately fair and project 199-16 (PSZ2G081) could not be observed because of its attenuation limit of 0.75. Project 169-22 GJ526North (1 rep. 1hr) was observed in less restrictive conditions (attenuation limit of 0.65).

09:00-11:00; Angel: tau_225GHz about 0.4 with an expected increase. Switching to EMIR, also anticipating the scheduled session of POLAMI. The second block of short-spacings of EMIR backup program 124-22 was observed until 11:05.

11:05-17:00; Ioannis, Albrecht: POLAMI.

17:00-21:00; Stefano: I take over after the POLAMI session at 17:20 local time. The wind is rising. After a couple of calibrations (Neptune and Uranus, scans 184, 186), I start 105-22 with source B213-C6 (4 repetitions). The weather then degrades as the shift ends.

Mother Nature delights us today with a wonderful gift:

attachment:collage_tramonto_20230116_27_28_29_crop2_sm2.JPG attachment:IMG_6962_mod_sm.JPG

21:00-03:00; Damien: 23h; weather too unstable to start yet. No observations.

Jan. 17th, Tuesday

03:00-10:30; No observations. Dense freezing fog and wind. Deicing active.

DailyReportsNika2Pool100123 (last edited 2023-01-18 15:16:09 by NikaBolometer)